### v1

# Enhancing Redundancy-based Automated Program Repair by Fine-grained Pattern Mining

Abstract.
Redundancy-based automated program repair (APR), which generates patches by referencing existing source code, has gained much attention since they are effective in repairing real-world bugs with good interpretability. However, since existing approaches either demand the existence of multi-line similar code or randomly reference existing code, they can only repair a small number of bugs with many incorrect patches, hindering their wide application in practice. In this work, we aim to improve the effectiveness of redundancy-based APR by exploring more effective source code reuse methods for improving the number of correct patches and reducing incorrect patches. Specifically, we have proposed a new repair technique named Repatt, which incorporates a two-level pattern mining process for guiding effective patch generation (i.e., token and expression levels). We have conducted an extensive experiment on the widely-used Defects4J benchmark and compared Repatt with eight state-of-the-art APR approaches. The results show that our approach complements existing approaches by repairing 15 unique bugs compared with the latest deep learning-based methods and 19 unique bugs compared with traditional repair methods when providing the perfect fault localization. In addition, when the perfect fault localization is unknown in real practice, Repatt significantly outperforms the baseline approaches by achieving much higher patch precision, i.e., 83.8%. Moreover, we further proposed an effective patch ranking strategy for combining the strength of Repatt and the baseline methods. The result shows that it repairs 124 bugs when only considering the Top-1 patches and improves the best-performing repair method by repairing 39 more bugs. The results demonstrate the effectiveness of our approach for practical use.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; December 2023; price: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Automated program repair (APR) techniques have the potential to significantly reduce the debugging overhead of human developers and thus have been widely studied in the last decades. To date, a large number of APR approaches have been proposed and developed (Wen et al., 2018; Xiong et al., 2017; Xuan et al., 2017; Liu et al., 2019; Tan et al., 2016; Weimer et al., 2009; Long and Rinard, 2016; Jiang et al., 2018; Xin and Reiss, 2017b; Nguyen et al., 2013; B. Le et al., 2016). A typical APR approach takes a faulty program and a set of test cases, where at least one triggers the bug in the program, as inputs, and produces one or more (if possible) plausible patches that can make all the test cases pass. The repair process can be typically viewed as a search problem, where the search space is all possible programs. Therefore, the core challenge for APR techniques is how to effectively and efficiently locate the desired patches within limited computing resources and time budget. To facilitate this process, existing techniques have employed multiple data sources, e.g., similar code (Jiang et al., 2018; Xin and Reiss, 2017b), historical patches (Le et al., 2016, 2017), repair templates (Liu et al., 2019; Kim et al., 2013; Saha et al., 2017), etc., and leveraged diverse search algorithms, e.g., random search (Qi et al., 2014), genetic programming (Weimer et al., 2009; Le Goues et al., 2012), and advanced deep learning (Zhu et al., 2021; Ye et al., 2022b; Jiang et al., 2023a, 2021), aiming to effectively refine the search space and thus improve the repair ability and efficiency.
Among existing APR techniques, redundancy-based (Chen, 2018; White et al., 2019; Yang, 2021; Chen and Monperrus, 2018) techniques have attracted much attention and have gained big success. These techniques are on the basis of the plastic surgery hypothesis that “changes to a codebase contain snippets that already exist in the codebase at the time of the change, and these snippets can be efficiently found and exploited (Barr et al., 2014)”. Therefore, redundancy-based APR approaches generate patches by searching and referencing existing code. For example, GenProg (Weimer et al., 2009; Le Goues et al., 2012), the pioneer of modern APR techniques, generates patches by directly reusing existing code under the guidance of genetic programming algorithms. Although the recent advance of large-scale pre-trained models has shown the promise to repair more bugs (Zhu et al., 2021; Jiang et al., 2023a; Ye et al., 2022b), they still suffer from the interpretability issue due to the complexity and the random nature of deep neural networks, and mainly repair simple bugs (Ye et al., 2022b). As reported by previous studies (Liu et al., 2020; Motwani et al., 2022; Wang et al., 2019), the patches generated by SimFix, a representative redundancy-based APR approach, are of high quality under various testing scenarios. In addition, redundancy-based techniques also complement those learning-based techniques (Zhu et al., 2021; Ye et al., 2022b; Jiang et al., 2023a) (Our evaluation results in Section 5 further prove it). Additionally, redundancy-based techniques tend to have better interpretability as they are on the basis of the plastic surgery hypothesis, and the patches are typically generated by referring human-written code snippets. In particular, they can repair a portion of relatively complex bugs (Yang et al., 2023; Jiang et al., 2018; Xin and Reiss, 2017b).
Although existing redundancy-based techniques are demonstrated to be effective for repairing real-world bugs, they still face two major challenges: (1) How to effectively and efficiently locate the reference code elements among a large-scale codebase? Existing approaches identify similar code snippets as the references for patch generation by leveraging code features like structures (Jiang et al., 2018; Yang et al., 2023) and token sequences (Xin and Reiss, 2017b). The idea behind them is that similar code snippets have a higher possibility to implement the same functionality and thus can guide bug fixes. However, as it will be shown in our preliminary study (Section 2.1), the coarse-grained similar code snippets for patch generation are indeed rare in real practice. The limited number of similar code snippets hinders them from repairing a large number of bugs and thus decreases their usability. (2) How to reuse existing code for constructing new patches? Existing approaches reuse code by comparing the difference between faulty and reference code via abstract syntax trees (Jiang et al., 2018; Xin and Reiss, 2017b) or program dependency graphs (Yang et al., 2023). However, code elements under different contexts tend to vary greatly in structures (Rattan et al., 2013), limiting the effectiveness of existing techniques, i.e., repairing a small number of bugs with many incorrect patches.
As explained above, existing approaches reuse code either completely randomly (e.g., RSRepair (Qi et al., 2014) and GenProg (Weimer et al., 2009)) or depending on the existence of similar code snippets of multiple lines (e.g., SimFix (Jiang et al., 2018) and TransplantFix (Yang et al., 2023)). As it will be presented in our preliminary study, though the reusable code elements for patch generation may appear at different granularities, most of them actually exist at the finer-grained level, e.g., 89.3% reusable code elements contain less than three tokens (Section 2.1). In this paper, we investigate the possibility of improving redundancy-based APR techniques by reusing code elements at a finer-grained level. Specifically, the basic idea of our approach is that fine-grained code elements related to similar code semantics tend to co-appear nearby in the program, which has been well-studied by existing research (Hindle et al., 2012; Ray et al., 2016; Khanfir et al., 2022; Tu et al., 2014). We call this phenomenon locality property of source code. Based on this property, we proposed a new APR approach that effectively guides patch generation by identifying the usage patterns of fine-grained code elements for confining the patch space and thus overcomes the first challenge of redundancy-based APR. To overcome the second challenge, i.e., reusing existing code under different contexts, we designed a new code representation method, i.e., S-TAC, which decomposes complex code expressions and statements into a unified form by ignoring context-specific features (e.g., code structures and operators). It can reduce the negative impact of code structures on code matching during patch generation but still preserve the locality property of source code.
To evaluate the performance of our approach, we have implemented it as an APR tool named Repatt and conducted an extensive study by comparing it with eight state-of-the-art APR approaches, including three best-performing traditional repair techniques and five latest deep learning-based techniques. The experimental results show that our approach complements existing approaches by correctly repairing many unique bugs. Specifically, Repatt repairs 19 and 15 unique bugs that cannot be repaired by the traditional and deep learning-based methods, respectively. In particular, 9 bugs repaired by Repatt have never been repaired by all the baselines. Additionally, when the perfect fault localization is unknown, which is a more realistic scenario, Repatt can significantly outperform the baseline approaches by achieving 33.8%-51.7% higher patch precision. Additionally, we also made the first attempt to combine the strength of Repatt and multiple existing approaches by further designing a patch ranking strategy. The evaluation result shows that the combined method can effectively repair 124 bugs when only considering Top-1 patches, 39 more bugs than the best-performing method. Our results demonstrate that our approach is indeed effective in improving the performance of existing APR techniques.
In summary, we make the following major contributions.
  • A new redundancy-based APR technique that can flexibly reuse source code at finer-grained granularities.
  • A novel code representation method, which overcomes the diversity of code contexts, for better code search and reuse.
  • An extensive study to evaluate the performance of our approach by comparing it with state-of-the-art approaches.
  • The first attempt to combine the strength of multiple APR approaches, revealing its promise for practical use.
  • We make all our experimental results and implementations publicly available to facilitate replication and comparison.
The remaining of the paper is organized as follows. Section 2 introduces our preliminary study and running examples to motivate our approach. Section 3 introduces the details of our approach. Then, Section 4 presents the experimental setup while Section 5 analyzes the evaluation results. We discuss the threats to validity and related work in Section 6 and 7, respectively. Finally, Section 8 concludes the paper.
2. Motivation
2.1. Preliminary Study
Figure 1. The length distribution of code elements that can be found in the faulty program for repairing related bugs.
Although many redundancy-based APR techniques have been proposed and produced promising results (Jiang et al., 2018; Xin and Reiss, 2017b; Liu et al., 2020; Motwani et al., 2022; Wang et al., 2019), these techniques reuse existing code elements by either random search without considering the relevance of reference code (Weimer et al., 2009; Le Goues et al., 2012; Weimer et al., 2013; Qi et al., 2014) or searching similar code at a coarse-grained granularity, e.g., code blocks (Jiang et al., 2018; Xin and Reiss, 2017b). However, no study explores the potential of searching and reusing finer-grained code elements, e.g., tokens, for patch generation. To investigate the feasibility and necessity of reusing finer-grained code elements for APR, we have conducted a preliminary study by following the idea of Barr et al. (Barr et al., 2014). Specifically, the objective of our study is to understand the granularity of reusable code elements for patch generation. To achieve that, we performed a greedy search process. When given a bug and its associated patch, we first extracted the newly added code elements in the patch and searched for them in the faulty program from coarse-grained to finer-grained. That is, if a code element (e.g., “a + f(b,c)”) was not found, we would further decompose it into finer-grained code elements based on the syntax tree structure of the code (e.g., “a”, “+”, and “f(b,c)”). This process was iteratively proceeded until the elements cannot be further decomposed into finer-grained ones (e.g., variables). Finally, we analyzed the granularity of code elements that can be found in faulty programs. In particular, we use the number of tokens to measure the granularity of code elements since code structures in the Abstract Syntax Tree (AST) cannot explicitly reveal the granularity of code elements. For example, an expression in the AST can represent either a simple variable like “a” or a complex expression like “f(a, b)”. We conducted our empirical study on two widely-used datasets of real-world bugs, i.e., Defects4J (Just et al., 2014) and Bugs.jar (Saha et al., 2018). In particular, we filtered bugs whose patches involve more than one Java file since they are still hard to repair at present (White et al., 2019; Jiang et al., 2018; Xiong et al., 2017; Xia et al., 2023; Jiang et al., 2023a; Zhu et al., 2021; Chen et al., 2019; Li et al., 2020). As a result, 564 bugs in Defects4J and 609 bugs in Bugs.jar from 24 diverse projects were included. Figure 1 presents the statistical results. The x-axis denotes the number of tokens in a code element and the y-axis denotes the percentages of code elements that can be reused for patch generation. Please note that a patch may involve multiple code elements. The results show on average more than 89.3% code elements contain less than three tokens, and 42% elements only contain one token. In contrast, only a very limited portion of code elements exist at a coarse-grained granularity (i.e., containing more than three tokens). In particular, this is not a special case but holds over different programs in our study. The result indicates that the reusable code elements indeed exist at a fine-grained granularity, which will significantly limit the effectiveness of existing redundancy-based APR approaches. Moreover, it also reflects that designing APR approaches by reusing finer-grained code elements is promising. However, achieving accurate and fine-grained code reuse for patch generation is challenging. The reasons are twofold: (1) It will significantly enlarge the search space of candidate patches, and thus make it harder to correctly and efficiently find the desired code elements. (2) Plausible (i.e., can pass all the test cases) but incorrect patches will be more easily generated due to the problem of weak test suite (Qi et al., 2015; Xin and Reiss, 2017a; Xiong et al., 2018). It will cause heavier manual effort for patch validation and thus significantly affect the usability of the APR approach in practice. To overcome this challenge, in this work, we propose an effective APR approach that can efficiently locate the desired code elements for patch generation. The basic idea of our approach is that code elements related to similar code semantics tend to co-appear nearby in the programs. For example, the variables “numerator” and “denominator” tend to be used together. We call this phenomenon the locality property of source code. Consequently, it would be possible to employ those code elements, which tend to co-appear with the ones that already exist in the faulty location, for patch generation. As it will be presented in our evaluation, the locality property does help to repair real-world bugs. The details will be introduced in Section 3.
2.2. Running Example
Motivated by our preliminary study, reusing finer-grained code elements for patch generation and boosting the effectiveness of redundancy-based program repair is feasible and promising. In this section, we employ running examples of real-world bugs to further explain how the locality property of source code can be leveraged to guide patch generation. Listing 1 presents the patch code of Codec-3 from the Defects4J (Just et al., 2014), where a line starting with “+” denotes newly added code while starting with “-” denotes code to be deleted.
452   if ((contains(value, 0 ,4, "VAN ", "VON ") || ...)){
453       //-- obvious germanic --//;
454       result.append(’K’);
455 - } else if(contains(value, index + 1, 4, "IER")) {
+ } else if(contains(value, index + 1, 3, "IER")) {
456       result.append(’J’);
457  } else {...
Listing 1: Patch code of Codec-3 in Defects4J
In this example, the constant parameter “4” was mistakenly used at line 455, and the desired parameter is “3”. This bug cannot be fixed by existing redundancy-based APR techniques due to either the large search space (e.g., GenProg (Weimer et al., 2009; Le Goues et al., 2012), RSRepair (Qi et al., 2014), etc.) or the inexistence of multi-line similar code snippets for reference (e.g., SimFix (Jiang et al., 2018), transplantFix (Yang et al., 2023), etc.). However, when it comes to the fine-grained token level, we will find that the constant number “3” usually co-appears with the string “IER” elsewhere in the program since the former really denotes the length of the latter string. In other words, these two tokens are strongly correlated to each other (i.e., the locality property of source code). As a result, it will be feasible and easier to effectively find the desired code elements for patch generation under the guidance of this property. In fact, the locality property of source code may not only exist at the token level but may also exist at higher levels, e.g., expression or statement levels. For example, Listing 2 presents the patch code of Gson-17 from the Defects4J benchmark, while Listing 3 presents a reference code that can be used for patch generation. According to the reference code, the expression “in.peek()” may be correlated with the expressions “JsonToken.NULL”, “in.nextNull()” and “return null”. In other words, the reference code can provide the required code elements for generating the desired patch in this example. However, it also can be seen that although these expressions can be taken for patch generation, how to utilize them is still challenging since the possible combinations of code changes according to the reference code are still too many, such as replacing the throw statement with “in.nextNull()” or inserting “in.nextNull”. In addition, replacing the conditional expression also does not fix the bug due to the incorrect operator “!=”. In particular, the code structures of reference code and the faulty code may also be different in practice, making patch generation harder.
98    public Date read(JsonReader in) throws IOException {
99 -  if (in.peek() != JsonToken.STRING) {
100 -     throw new JsonParseException("The date ...");
+  if (in.peek() == JsonToken.NULL) {
+     in.nextNull();
+     return null;
101     }
102     Date date = deserializeToDate(in.nextString()); ...
Listing 2: Patch code of Gson-17 in Defects4J
1  if (in.peek() == JsonToken.NULL) {
2      in.nextNull();
3      return null;
4  }
5  in.beginArray();
Listing 3: The reference code for repairing Gson-17
To address this challenge, we propose a novel code representation method that decomposes complex expressions into a unified simple form, which ignores operators and complex code structures, e.g., !=, for, while, etc., but still preserves the locality property of source code. For example, the transformed representation of the faulty code and the reference code shown in Listings 2 and 3 are presented in Figure 4. As a result, we can generate candidate patches by aligning these two representations.
3. Framework
Figure 2. Overview of our approach Repatt
In this section, we introduce the details of our approach (named Repatt). Figure 2 presents the overview of Repatt. In general, Repatt generates candidate patches by mining the usage patterns of code elements at different granularities as guidance based on the plastic hypothesis and the locality property of source code. Particularly, in order to repair bugs that demand code changes at different granularities, Repatt comprises two components: token-level pattern mining and expression-level code search. The former aims to repair bugs that demand code changes within an individual line (ref. Listing 1), while the latter aims to repair bugs that demand code changes across multiple code lines (ref. Listing 2).
Specifically, the token-level pattern mining is an offline process whose target is to construct a code pattern database that can be effectively and efficiently queried in the subsequent online repair process. On the contrary, the expression-level code search is an online process that aims to search for a set of code elements that can be referenced for patch generation. The reasons for their different search strategies are twofold: (1) As aforementioned, token-level pattern mining aims to repair single-line bugs. Therefore, we can only consider the usage patterns of tokens co-appearing within single lines, which can be small-scale. In contrast, the expression-level code usage patterns will tend to produce a much larger search space when considering all possible combinations of expressions from different lines of code. (2) Token-level patterns tend to produce many small code changes and thus may severely affect the efficiency of the repair process, while the expression-level patterns correlated to a certain location (i.e., the faulty code) will be very limited according to the statistical results in our study (refer to Figure 1). Therefore, we confine the token-level patterns to those that are frequent. In other words, we need to analyze whether the two patterns are the same or not for tokens.
3.1. Offline Pattern Mining
Input: ESs: sequences of code element embeddings (IDs)
Output: trees: frequent pattern in the prefix trees form.
1 Function build(ESs):
[…]
             buildTree(S, trees, updated)   // Build trees per each line
3
      return trees   // Embedding trees based on the complete program
4
5Function buildTree(seq, trees, updated):
6       start $\leftarrow$ 0 while start $<$ len(seq) do
             root $\leftarrow$ findOrCreate(trees, seq[start])   // Query existing trees
             root.sup $\leftarrow$ root.sup + 1   // Increase the token frequency
[…]
8Function traverse(tree, seq, skip, length, updated):
9       if length $\geq$ MAX_LEN or len(seq) ==absent$=$ = 0 then
10             return
11       else if skip $<$ MAX_SKIP then
             traverse(tree, seq[1:], skip+1, length)   // Skip the token
12
      node $\leftarrow$ tree.childNode(seq[0])   // Do not skip the token
13       if node is None then
14             node $\leftarrow$ create a new node of id seq[0] tree.setChildNode(seq[0], node)
15       if node is not in updated then
             node.sup $\leftarrow$ node.sup + 1   // Increase token frequency
[…]
17       traverse(node, seq[1:], skip, length + 1)
Algorithm 1 Tree Building and Pattern Construction
According to the preliminary study presented in Section 2.1, most of the reusable code elements for patch generation exist at very fine-grained granularity, i.e., one or two tokens. Therefore, to reuse code elements at the token level, we proposed a new code pattern mining algorithm in Repatt by leveraging the locality property of source code. Specifically, Repatt performs an offline pattern mining process over the given faulty project, which aims to construct a database of token usage patterns that can be used for facilitating the online patch generation process in the subsequent repair process. As aforementioned, we only consider the token patterns within single lines to confine the search space and avoid involving too much noise. In particular, it is common that the tokens used in different code lines share both commonalities and diversities. For example, the token sequence in some code lines can be (“a”, “b”, “c”), while it may also be (“a”, “d”, “c”) in some other code lines. As a result, the sequence (“a”, “c”) is commonly used in both code lines (i.e., commonality) while the tokens “b” and “d” are specific to certain code contexts (i.e., diversity). To preserve the semantics of programs, we keep the token orders and design a skip-fashion pattern mining algorithm to address the problem of diversity among code lines. Also, Repatt embeds each token with a unique id for facilitating the token comparing process. In particular, Repatt removes the separators (e.g., “,” and “(”) and structure keywords (e.g., “if” and “for”) since they are unhelpful for token-level patch generation.
Specifically, to improve the pattern mining process, we design a data structure – Prefix Embedding Tree, which is defined as follows.
Definition 3.1 ().
(Prefix Embedding Tree): a prefix embedding tree constitutes a set of nodes, each of which is a tuple of $t=\left\langle tok,id,p,C,sup\right\rangle$, where $id$ denotes a unique embedding of a certain token $tok$, $p$ denotes the parent node of $t$, $C$ denotes a set of child nodes of $t$, while $sup$ represents the frequency of the token sequence from the root to the current node $t$.
Furthermore, we use $t.childNode(id)$ to obtain the child node of $t$ with the embedding $id$, and use $t.setChildNode(id,chd)$ to set the node $chd$ as the child node of $t$. According to these notions, we present the pattern mining process in Algorithm 1. In general, given the faulty program, Repatt first decomposes all the code lines in the program into a set of token sequences (i.e., ESs), where each token sequence corresponds to one line of code. Then, it constructs the prefix embedding trees (i.e., trees), which stores the token patterns with corresponding usage frequencies.
Specifically, for each token sequence $S\in$ESs (line 5), Repatt either creates new trees using it or expands existing trees by increasing the frequencies of corresponding patterns (lines 10-11 and 20-26). In particular, we use the function findOrCreate(trees, tok) to find the tree in trees, whose root node corresponds to the token tok, or create a new tree in case it does not exist. Additionally, Repatt employs a skip-fashion pattern mining strategy in this process. That is, for a given token, Repatt will consider both including (lines 20-27) and skipping (lines 18-19) it in the constructed patterns. In particular, we set two hyper-parameters MAX_LEN (line 16) and MAX_SKIP (line 18) to control the max number of tokens in a pattern and the max number of tokens that are allowed to skip. In our evaluation, we will also study their effects on our approach.
Taking the faulty code Codec-3 (Listing 1) as an example, the token sequence of the faulty code line will be $S$=(“contains”, “value”, “index”, “+”, “1”, “4”, “IER”), based on which the constructed prefix embedding trees are presented in Figure 3, where we use the red line to highlight the patterns related to the token sequence $S$. Consequently, the token sequence from the root node to the leaf node denotes a possible pattern and the number denotes its frequency. For example, the frequency of the pattern (“contains”, “value”, “1”, “IER”) is 3, while the pattern (“contains”, “value”, “1”, “3”) is 5. In particular, each distinct token will be the root node of an individual prefix embedding tree. For instance, we present two prefix embedding trees in Figure 3 that are related to the token sequence $S$. As a result, when given a buggy sequence, we can only search a very limited number of embedding trees whose root tokens are included in the sequence, which can promote the searching process. In our pattern mining process, given the minimal frequency of patterns as MIN_SUPPORT, the patterns whose frequency is smaller than MIN_SUPPORT will be deleted. A higher frequency represents a higher possibility of the pattern to be common and reusable for program repair. In the repair process, given the faulty code token sequence $S^{\prime}$, Repatt will search all possible patterns over the embedding trees and generate patches by referencing them.
Figure 3. Partial prefix embedding trees. We use concrete tokens to represent the token ids for ease of understanding.
3.2. Online Code Search and Representation
The expression-level code search aims to find code usage patterns across multi-lines of code for reference, which may cause a large search space if it is also considered like the token-level pattern mining since there may be hundreds or thousands of lines of code even in a single method. As a result, the expression-level pattern mining process is designed as an online code search process, which automatically identifies a set of reference code when given the faulty code. Following existing work (Jiang et al., 2018), Repatt extracts no more than three lines of code respectively before and after the given faulty line as the faulty snippet, and then finds the $N$ most similar snippets via measuring the code similarity between the faulty and reference snippets. Specifically, Repatt extracts a vector from each snippet, where each element in the vector represents a feature and the feature value represents the number of corresponding AST node types in the snippet by following the study (Jiang et al., 2018). Then, we calculate the cosine similarity (Salton, 1988) between two vectors for code ranking.
After obtaining the reference code, existing redundancy-based APR approaches usually try to generate patches by directly referencing the similar code snippets (Yang et al., 2023; Jiang et al., 2018), which is valid if the reference snippets are very similar to the faulty one. However, the contexts of desired reference snippets can be diverse and largely different from the faulty code in real practice, where existing methods will be less effective. To address this issue, we propose a new code representation method, named “Simplified Three-Address Code” (abbr. S-TAC). It unifies the code representations by ignoring the code structures under different contexts but still preserves the locality property of source code.
Definition 3.2 ().
(Simplified Three-Address Code): an S-TAC is a triple of $\left\langle T,t_{1},t_{2}\right\rangle$, where $T$ is an intermediate symbol to represent the expression correlated to $t_{1}$ and $t_{2}$, while $t_{1}$ and $t_{2}$ are two expressions that are either the intermediate symbols or simple items of variables, literal values, types, and function calls. In particular, $t_{1}$ and $t_{2}$ can be null if they are keywords, e.g., break and return.
Additionally, we use $T:=t_{1},t_{2}$ to represent $\left\langle T,t_{1},t_{2}\right\rangle$ when there is no ambiguity. Compared with traditional Three-Address Code (TAC) (Aho et al., 2020), S-TAC ignores both the coarse-grained code structure information, e.g., if, and the fine-grained operators, e.g., “+”. The reason for ignoring the structure information is to make the source code from different contexts applicable for patch generation. For example, the conditional expression $a>b$ in if statements can match that either in for statements or in trinary conditional expressions, enabling a more flexible code match. Similarly, the fine-grained operators may also be specific to certain contexts and thus affect the matching and reuse of code. In other words, our S-TAC representation can better reflect the locality property of source code while reducing the noise induced by different contexts.
Taking the code shown in Listing 2 and 3 as examples, the S-TAC form for the if condition from the faulty code is $T_{1}:=in,peek()$, $T_{2}:=T_{1},JsonToken.STRING$, while it will be $T_{1}:=in,peek()$, $T_{2}:=T_{1},JsonToken.NULL$ for the reference code, where $T_{1}$ and $T_{2}$ are intermediate symbols. In this way, the two conditions can be better matched with each other. We will also evaluate the contribution of our S-TAC form in the experiment (Section 5).
3.3. Reference Code Adaptation
When given the faulty code, we can obtain a set of reference code based on the previous two processes. Then for each reference code pattern, Repatt tries to generate candidate patches by matching the faulty code to the reference code. Algorithm 2 presents the details of the matching process. In general, Repatt performs a greedy match between the faulty code and the reference code. Then, it generates patches according to the difference between them. In other words, given the token (or S-TAC) sequences for both faulty (i.e., BS) and reference code (i.e., RS), the matching algorithm will return a set of pairs (i.e., PS), based on which Repatt generates candidate patches according to a set of pre-defined rules.
[…]
Input: BS: a sequence of tokens or S-TACs that represent the bug. RS: a sequence of tokens or S-TACs that represent the reference code.
Output: PS: a set of original and target node pairs.
1 Function matchElement(BS, RS):
       PS [$\leftarrow$[]← [ ], bf $\leftarrow-1$, rf $\leftarrow-1$ cs $\leftarrow LCS(BS,RS)$   // The longest common sequence
2       if  cs.size() $>$ 0 then
3             foreach c in cs do
[…]
4                   bf $\leftarrow BS.indexOf(c.fst)$ rf $\leftarrow RS.indexOf(c.snd)$
[…]
5
6       PS normal-$\leftarrow$ PS $\cup$ tryToMatchParent(PS) return PS
7Function tryToMatchParent(PS):
8       result [$\leftarrow$[]← [ ] foreach $\left\langle a,b\right\rangle$ in PS do
9             parent $\leftarrow a.getParent()$ children $\leftarrow parent.getChildren()$ if PS.FirstSet().containsAll(children) then
10                   ts [$\leftarrow$[]← [ ] foreach c in children do
11                         ts.append(PS.getSecond(c).getParent())
[…]
13
14      return result
Algorithm 2 Reference Code Matching
Figure 4. Matching results of partial S-TACs in Gson-17 shown in Listing 2. $b_{i}$ represents the S-TAC form of the faulty if statement while the $p_{i}$ corresponds to the if statement in the reference code.
Finally, according to the constructed mapping (i.e., PS), Repatt will generate candidate patches from each pair $(a,b)\in PS$ according to the following rules (we use $a$ to represent the associated AST node in what follows for ease of presentation): (1) replace $a$ with $b$ if the value type of $b$ is compatible with that of $a$; (2) insert $b$ before and after $a$ if $b$ is a standalone statement, e.g, expression statement; (3) insert $b$ before $a$ if $b$ is a conditional expression. In particular, we disable delete operations for patch generation since they tend to generate incorrect patches (Tan et al., 2016; Jiang et al., 2018). Moreover, during this process, Repatt will automatically check the validity of the new code $b$ via static analysis for efficiency, such as whether the variables used by $b$ are all defined in the faulty location. Taking the mapping shown in Figure 4 as an example, one candidate modification will be replacing the expression in.peek()!=JsonToken.STRNG with in.peek()==JsonToken.NULL according to the mapping $(b2,p2)$.
3.4. Patch Ranking
After generating candidate patches, they will be evaluated against the test suites associated with the faulty project. To make the correct patches be evaluated as early as possible and reduce incorrect patches, we have proposed a hierarchical patch ranking strategy. As demonstrated in our preliminary study, most of the reusable code elements are at the fine-grained granularity. Furthermore, previous studies (Mechtaev et al., 2015) also show that correct patches tend to involve small code changes. Therefore, Repatt ranks all the patches generated by the finer-grained token-level code changes higher than those generated by the coarse-grained expression-level code changes. In other words, the patches generated by token-level code changes will be evaluated first and then those by expression-level code changes. Please note that the online pattern mining process introduced in Section 3.1 can be proceeded independently ahead of time and the constructed patterns can be reused for repairing different bugs on demand.
Then, regarding the patches generated by expression-level code changes, Repatt simply takes the similarity in code search (refer to Section 3.2) as the ranking score by following previous studies (Jiang et al., 2018). In contrast, for patches generated by token-level code changes, we consider two important factors: the frequency of the reference code pattern and the similarity between the faulty and fixed code by following existing studies (Mechtaev et al., 2015; Jiang et al., 2018). The ranking score is defined by Formula 1,where $freq$ denotes the frequency of a pattern and $max\_freq$ denotes the max $freq$ of patterns referenced for generating patches. $L_{orig}$ and $L_{fixed}$ represent the length of source code (i.e., number of tokens) before and after applying a patch, while LevenshteinDist represents the Levenshtein distance (Levenshtein, 1966) between the faulty and the fixed code. As a result, a patch with a higher score will rank higher and thus will be validated earlier by running the test cases in the faulty project.
(1) $\small score=0.5*\frac{freq}{max\_freq}+0.5*(1-\frac{{LevenshteinDist}}{max({L_{orig}},{L_{fixed}})})$
4. Experiment Configuration
We address the following research questions in the study.
  • RQ1: How effective is our approach in repairing real-world bugs?
  • RQ2: Can our approach improve the state-of-the-art APRs?
  • RQ3: What is the contribution of each component in Repatt?
  • RQ4: What is the impact of the pattern frequency in Repatt?
4.1. Subjects and Baselines
In our evaluation, we employed the widely-used Defects4J benchmark (Just et al., 2014) by following existing studies (Xin and Reiss, 2017b; Xuan et al., 2017; Xiong et al., 2017; Yang et al., 2023; Jiang et al., 2023a; Zhu et al., 2021; Jiang et al., 2018; Liu et al., 2019). In particular, we used both Defects4J v1.2 and v2.0 to demonstrate the generality of our approach. Specifically, v1.2 includes 395 bugs from six large-scale real-world projects and v2.0 includes 440 additional bugs from 12 real-world projects.
Furthermore, to show the effectiveness of Repatt, we compared it with eight state-of-the-art APR approaches from different categories, i.e., SimFix (Jiang et al., 2018), TransplantFix (Yang et al., 2023), TBar (Liu et al., 2019), Recoder (Zhu et al., 2021), SelfRepair (Ye et al., 2022b), CodeT5 (Jiang et al., 2023a), CodeGen (Jiang et al., 2023a) and PLBART (Jiang et al., 2023a). Specifically, SimFix and TransplantFix are the two latest and representative redundancy-based methods that repair bugs by referencing similar code. TBar is the best-performing template-based repair technique that generates candidate patches by a set of manually defined patch templates. Finally, Recoder and SelfRepair are the latest and best-performing deep learning methods that are specially designed for program repair, while CodeT5, CodeGen, and PLBART represent the most recent advance in APR techniques by adopting large-scale pre-trained language models. These baselines are the best-performing APR approaches using different technologies and their complete experimental results are available. By comparing our approach with these diverse baselines, we would like to analyze the overall effectiveness of our approach from different perspectives and make the conclusions reliable.
4.2. Configuration and Metrics
For each bug, Repatt first constructs the token-level code patterns based on the complete faulty program under repair before the online repair (see Section 3.1). The construction is actually efficient and on average took about three minutes in our experiment. Then, given the faulty line of code, Repatt first generates at most 200 patches based on the constructed token-level code patterns and then generates at most 1000 patches based on the expression-level code patterns due to its large search space. Finally, all the patches will be ranked based on the ranking strategy introduced in Section 3.4. Additionally, we assign each bug 90 minutes for repair (Le et al., 2016; Wen et al., 2018; Saha et al., 2017). We empirically set the default value of MIN_SUPPORT in token-level pattern mining as 3, and the MAX_SKIP and MAX_LEN as 8 and 5, respectively. For baselines, we adopt their published experimental results directly in their open-source repositories. In particular, for the three large-scale pre-trained deep learning-based methods, we adopt the fine-tuned models CodeT5-large, CodeGen-6B, and PLBART-large, which achieved the best repair results (Jiang et al., 2023a).
In our experiment, Repatt generates at most 3 candidate patches that can pass all the test cases (i.e., plausible patches) for each bug since some baseline approaches did not report the ranking of patches, e.g., SelfRepair. Following previous study (Xin and Reiss, 2017b; Jiang et al., 2018; Liu et al., 2019; Zhu et al., 2021), a patch is deemed to be correct iff it is semantically-equivalent to the developer patch by manual check. Finally, we report the number of correctly repaired bugs (equivalent to the well-known recall) and the precision of patches (the ratio of correctly repaired bugs to all the bugs that have plausible patches). Please note the recall and precision here are different from those in the deep learning field since all the bugs are expected to be repaired, i.e., the true negative in deep learning field does not make sense here. All experiments were conducted on a server with Ubuntu 18.04, equipped with 128GB RAM and a processor of Intel(R) Xeon(R) E5-2640.
5. Result Analysis
Table 1. Number of bugs repaired by different methods with perfect fault localization.
Project Recoder TBar SimFix TransplantFix SelfRepair CodeT5 CodeGen PLBART Repatt Combine
Defects4J v1.2 (395 bugs) Mockito 2/- 3/3 0/0 3/3 3/3 4/5 3/4 2/3 1/2 2/5
Closure 23/- 16/24 5/6 10/19 19/23 6/9 9/12 4/10 10/12 20/31
Chart 10/- 11/13 4/8 6/10 7/10 4/4 7/7 4/4 6/7 11/21
Lang 10/- 13/18 8/13 4/10 10/13 6/8 5/5 4/4 6/11 16/27
Math 18/- 22/35 14/26 12/25 21/25 12/14 12/14 13/16 15/25 22/51
Time 3/- 3/6 1/1 1/2 3/3 1/1 2/2 2/2 2/3 4/7
Defects4J v2.0 (440 bugs) Closure 0/- 0/0 1/1 0/3 1/1 1/1 1/1 1/1 1/1 2/5
Cli 3/- 1/7 0/1 4/8 8/9 2/3 4/5 2/3 4/9 4/15
Codec 2/- 3/6 1/1 2/4 8/9 1/1 2/3 1/1 4/6 3/8
Collections 0/- 0/0 0/0 0/0 1/1 0/0 1/1 0/0 0/0 0/0
Compress 3/- 2/12 0/4 4/10 6/8 3/5 3/6 3/5 4/13 7/18
Csv 3/- 2/6 0/2 1/2 1/1 1/2 2/3 2/4 1/3 2/7
Gson 0/- 1/4 2/2 1/2 1/1 0/0 0/0 0/0 2/3 2/5
JacksonCore 0/- 0/5 0/2 1/6 3/3 3/3 2/3 2/2 3/7 3/9
JacksonDatabind 0/- 2/17 1/10 7/20 8/10 4/7 3/6 2/4 10/18 11/33
JacksonXml - 0/1 0/0 0/0 1/1 0/0 0/0 0/0 0/0 1/1
Jsoup 7/- 6/17 1/2 3/8 6/8 4/6 5/7 4/5 5/13 10/25
JxPath 0/- 1/6 0/0 2/7 1/1 0/0 0/1 0/1 1/5 4/12
Total 84/- 85/180 38/79 61/139 108/130 52/ 69 61/80 46/ 65 75/138 124/280
Precision(%) - 47.2 48.1 43.9 83.1 75.4 76.3 70.8 54.3 44.3
5.1. Overall Effectiveness of Repatt (RQ1)
5.1.1. Perfect fault localization
As explained in Section 4.1, we evaluated the effectiveness of our approach by comparing it with eight state-of-the-art APR approaches. The repair results are presented in Table 1 when given the actual faulty locations by following existing studies (Jiang et al., 2023a; Yang et al., 2023; Zhu et al., 2021; Liu et al., 2019; Ye et al., 2022b). In each cell, X/Y denotes the repair method can generate at least one plausible patch (i.e., can pass all the test cases) for Y bugs on different projects, where X bugs are indeed repaired correctly. From the table we can see that our approach Repatt successfully repaired 75 bugs but also generated incorrect patches for other 63 bugs, leaving the patch precision 54.3%. In particular, 66 bugs were correctly repaired by the first plausible patch. As a result, developers may only consider the first patch generated by Repatt for manual validation to save human effort in practice. The results also show that our approach outperforms the state-of-the-art redundancy-based methods SimFix and TransplantFix by repairing 97.4% and 23.0% more bugs, indicating the effectiveness of our approach in reusing code at different granularities for patch generation. Please note that SelfRepair repaired 108 bugs when considering Top-50 patches. When considering the first (i.e., Top-1) plausible patches, it can only repair 34 bugs as reported in the original paper (Ye et al., 2022a). Please note that the latest advanced learning-based repair tool KNOD (Jiang et al., 2023b) did not report the detailed experimental results, i.e., the number of bugs that were correctly repaired by Top-1 patches. Nevertheless, even considering Top-10 patches, KNOD could only repair 20 and 13 bugs from Defects4J v1.2 and v2.0, respectively, still underperfoming our approach. Compared with the others, our approach can also achieve similar effectiveness. For example, Repatt achieved higher patch precision than TBar although repaired fewer bugs. In contrast, Repatt repaired more bugs than the pre-trained models (CodeT5, CodeGen, and PLBART) but achieved a lower patch precision. Moreover, as it will be presented in Section 5.1.2, our approach complements existing APRs by repairing many new bugs. Nevertheless, although our approach can effectively repair many bugs, there are still many bugs that cannot be repaired by Repatt because the patches cannot be assembled from individual reference code. To further improve the correct repair, enlarging the patch space may help, such as combining multiple reference code or searching a code base beyond the faulty project for patch generation. We plan to further investigate its performance in the near future.
5.1.2. Degree of complementary
According to the results shown in Table 1, although Repatt can successfully repair many bugs, it is unknown whether Repatt is still valuable in the face of many state-of-the-art APR approaches. Therefore, we further investigated the extent to which our approach complements the state-of-the-art baselines. We analyzed the overlaps of bugs repaired by both Repatt and the baselines. Specifically, we classified the baselines into traditional and deep-learning methods according to the patch generation techniques, and then compared Repatt with these two methods separately. Figure 5 presents the results. In summary, Repatt successfully repaired 19 unique bugs that the traditional APRs cannot fix, and 15 unique bugs that the deep learning methods cannot fix. In particular, when compared with all the eight baselines, our approach still can repair 9 unique bugs. For example, to repair the bug shown in Listing 1, the constant value “4” should be replaced by “3”. Without the guidance of the fine-grained reference code, it is nearly impossible for existing APRs to get the correct patch due to the large search space. In general, compared with traditional APRs, Repatt can more effectively utilize fine-grained reusable code elements for patch generation, while compared with deep learning methods, Repatt only depends on a small number of reusable samples rather than the large-scale training data, which enables Repatt to repair infrequent bugs. In conclusion, the results reflect that our approach indeed complements existing approaches.
(a) Traditional methods
(b) DL methods
Figure 5. Overlaps of bugs repaired by different approaches.
5.1.3. Automated fault localization
In addition, some of the baseline approaches (i.e., SimFix, TBar, TransplantFix, and Recoder.) were also evaluated in the scenario where the perfect fault localization was unknown in the original paper. Therefore, we also compared our approach with them when using the automated fault localization results by following existing research (Liu et al., 2019; Jiang et al., 2018; Zhu et al., 2021). Specifically, we utilized the commonly-used Ochiai(Abreu et al., 2007) algorithm that is implemented by the GZoltar toolset (Campos et al., 2012) to obtain a list of candidate faulty locations like existing APR tools (Liu et al., 2019; Jiang et al., 2018; Zhu et al., 2021). In our experiment, we use the default configuration of GZoltar. Table 2 summarizes the experimental results. From the table, we can see that the number of correctly repaired bugs drops sharply for all the repair approaches. For example, SimFix, TBar, TransplantFix, Recoder, and Repatt repaired 5.2%, 41.2%, 27.9%, 20% and 60% fewer bugs, respectively, compared with providing the perfect fault localization. The large decline in our approach is mainly due to the inaccurate fault localization results, which will cause a large search space since the finer-grained code reuse tends to generate candidate patches, making the correct patches unfound within the given time budget. Moreover, the patch precision of all the baseline approaches also incurred a large decline. For instance, the precision of SimFix, TBar, and TransplantFix drops 7.8%, 19.1%, and 26.9%, respectively. On the contrary, the precision of our approach was improved by 54.3%. The reason is that fine-grained code changes at the faulty location directly affect the results of failing tests and thus have a higher possibility to make them pass without affecting the results of passing tests, while fine-grained code changes to the correct code will have a larger impact on the passing tests. In contrast, the coarse-grained code changes by existing APRs tend to affect both failed and passed tests, and thus they will generate incorrect patches more easily under inaccurate fault localization due to the problem of weak tests (Qi et al., 2015; Xiong et al., 2018; Xin and Reiss, 2017a). Notice that the high precision of patches is very important since it affects the usability of APR techniques (Qi et al., 2015; Xiong et al., 2018), especially in the real-world repair scenarios where the prefect fault localization is unknown – high precision denotes less wasted human effort for manual validation.
Table 2. Repair result comparison when using SBFL
SimFix TBar TransplantFix Recoder Repatt Combine
Total 36/81 50/131 44/137 70/140 31/37 90/236
Precision (%) 44.4 38.2 32.1 50.0 83.8 38.1
5.2. Improvement over SOTA APRs (RQ2)
As presented in Section 5.1.2, Repatt complements existing APR approaches. In this research question, we aim to explore the possibility of our approach to improving existing APRs. Specifically, we combine Repatt with existing APRs by proposing a post patch ranking strategy and see whether it can further improve the best-performing APR. Given the patches generated by different APRs for a bug, we will rank the ones with smaller code changes higher by following the insights from both our study and the previous research (Mechtaev et al., 2015). In particular, we use GumTree (Falleri et al., 2014), a fine-grained and mature code differencing tool, to measure the change sizes of patches. If two patches have the same number of code changes, we use the patch precision to break the tie. In this study, we combined our approach with the three traditional APR approaches, i.e., Simfix, TBar, and TransplantFix. In this way, the combined method shares a similar repair pipeline to the individuals, i.e., not demanding large-scale training data. Therefore, the patch ranking will be Repatt$>$SimFix$>$TBar$>$TransplantFix based on their patch precision in Tables 1 and 2 for patches having the same change size.
(a) Using perfect FL
(b) Using SBFL
Figure 6. Source of correct patches in Combine
The experimental results are also presented in Tables 1 and 2 (i.e., Combine) when using the perfect fault localization and the SBFL, respectively. The results show that by combining the strength of different APR tools, Combine correctly repaired 124 bugs by the first patches, 39 more bugs compared with the best-performing APR (TBar for Top-1 patches), leaving the improvement as 45.9%. Similarly, when using the SBFL results, Combine also outperforms the best-performing Recoder by repairing 20 more bugs, and improves the individuals by at least 80% (vs Tbar). Our results show the promise to improve the repair capability of APRs by combining the strength of individuals. In particular, to analyze the contribution of our approach in the combination, we present the percentages of correct patches generated by each APR in Figure 6. About 41.1% and 25.6% correct patches were contributed by Repatt under different repair settings, indicating its large contribution to the overall effectiveness of the combined method.
However, though the combined method is effective in repairing much more bugs, the patch precision is still low, i.e., less than 45%. To improve the patch precision, existing patch filtering approaches (Tan et al., 2016; Xiong et al., 2018; Xin and Reiss, 2017a) can be further incorporated. As far as we are aware, this is the first attempt to combine the strength of different APRs. Although simple, the combined method already shows its promise. In the future, we plan to explore more effective combining methods and also incorporate the patch filtering process for improving the patch precision. Nevertheless, the results obtained from this study encourage further research in this direction.
5.3. Contribution of Each Component (RQ3)
Repatt incorporates two major components for patch generation, i.e., offline token-level pattern mining and expression-level code search. To evaluate the effectiveness of each component, we have repeated our experiment by removing each component, respectively. Table 3 presents the experimental results. The offline token-level pattern mining contributed 42 correct fixes and expression-level code search contributed 33 correct fixes, demonstrating that both components largely contributed to the overall effectiveness of Repatt. However, from the table, we can also observe that the expression-level code search tends to generate more plausible but incorrect patches compared with the token-level pattern mining, which is consistent with the conclusions in prior studies (Mechtaev et al., 2015). Our further analysis of the results shows that most of incorrect patches are generated due to referencing infrequent code. In contrast, token-level pattern mining takes the frequency into consideration, which potentially can effectively filter out many incorrect patches. In summary, the two components in Repatt complement each other.
In addition, as introduced in Section 3, our token-level pattern mining process incorporates a skip-fashion pattern mining process, which can find more potential reference patterns for patch generation. Therefore, we further conducted an experiment, where we only use the token-level repair but disable the skip-fashion in pattern mining (i.e., MAX_SKIP=0 in Algorithm 1). The column of “No Skip” in Table 3 presents the results. It shows that this skip-fashion mining process is effective since its removal caused 42-24=18 fewer bugs to be repaired. Moreover, the patch precision is also decreased from 65.6% to 44.4%, demonstrating the value of this process in Repatt. Additionally, we further evaluated the performance of the S-TAC form in our expression-level repair. Specifically, we did not transform the reference code into S-TAC. Instead, they will be decomposed into normal expressions without removing the code structures based on the abstract syntax tree. The experimental results are also presented in Table 3 (i.e., “No S-TAC”). After removing the S-TAC form, Repatt failed to repair any bugs by searching reference code online. We further analyzed the results and found that the contexts of most reference code from online search are different from the faulty code, making the code adaptation fail to generate correct patches. On the contrary, 22 incorrect patches were generated. In summary, the experimental results demonstrate that both the skip-fashion online pattern mining and the S-TAC form in the online search process are effective.
Table 3. Number of bugs fixed by each component.
Components Offline Mining Online Search No Skip No S-TAC
Total 42/64 33/85 24/54 0/22
Precision (%) 65.6 38.8 44.4 0
5.4. The Impact of Pattern Frequency (RQ4)
As introduced in Section 4.2, the threshold of pattern frequency is 3 by default. To investigate its impact on the effectiveness of our approach, we have conducted an additional experiment when using different frequency settings for MIN_SUPPORT, i.e., 2, 3, and 5. As a consequence, when adopting different settings, the performance of Repatt was not largely affected. Specifically, Repatt correctly repaired 43/91, 42/64, and 32/68 bugs by the token-level patterns when using the threshold as 2, 3, and 5, respectively. In other words, our approach is not very sensitive regarding the number of correct patches when the threshold is small (i.e., 2 or 3). However, when the threshold is too large (such as 5), Repatt tends to repair much fewer bugs. The reason is also clear as the large threshold will cause fewer patterns that can be used for patch generation. On the contrary, a smaller threshold (i.e., 2) tends to produce much more incorrect patches (i.e., 91-43=48) since the referenced patterns may be not general enough for reuse, which will affect the usability of Repatt. In summary, the recall and precision of program repair indeed contradict each other. Effective patch generation techniques are always in an urgent need. In this paper, we have moved forward towards this direction. In general, setting the threshold as 3 can produce relatively better results.
6. Discussion
Limitation: In this study, we have explored the feasibility of reusing finer-grained code elements for patch generation. For complex bugs that require multi-line changes, the reference code may not exist, where our approach will be less effective. In this case, our approach may potentially be combined with others since they complement each other according to our experimental results.
Threats: The threats to external validity mainly lie in the subjects used in our study. Since the Defects4J benchmark has been widely used by existing studies and contains more than 800 real-world bugs from 17 diverse projects, we believe the results can be reliable. However, the effectiveness of our approach on a broader range remains to be investigated. The threats to internal validity mainly lie in the implementation of our approach. To ensure the correctness and reliability of the experimental results, two authors have carefully checked the correctness of the implementation via code review. In addition, we also publish all our experimental results for replication and promoting future research.
7. Related Work
There are many automated program approaches have been proposed and achieved good results (Weimer et al., 2009; Long et al., 2017; Weimer et al., 2013; Pei et al., 2014; B. Le et al., 2016; Mechtaev et al., 2015; Nguyen et al., 2013; Mechtaev et al., 2018; Jiang et al., 2018; Xiong et al., 2017; Xuan et al., 2017; Le et al., 2016; Ghanbari and Zhang, 2018; Jiang et al., 2023a, 2021), among which redundancy-based APR approaches have been attracted much attention, such as GenProg (Weimer et al., 2009; Le Goues et al., 2012), AE (Weimer et al., 2013; Yuan and Banzhaf, 2018), RSRepair (Qi et al., 2014), ARJA (Yuan and Banzhaf, 2018), ssFix (Xin and Reiss, 2017b), SimFix (Jiang et al., 2018), CapGen (Wen et al., 2018), SearchRpepair (Ke et al., 2015), HERCULES (Saha et al., 2019) and so on. All these approaches are based on the “plastic surgery hypothesis” (Barr et al., 2014). The major difference among them lies in the adopted patch generation strategies. For example, GenProg reuses existing code randomly, ssFix searches similar code as references from a large codebase, while SimFix further incorporates history patches to refine the patch space. Similarly, GapGen considers the frequency of tokens for patch generation while HERCULES leverages the co-evolution of program elements. As explained in the introduction, our approach aims to improve the performance of redundancy-based APR by exploring finer-grained code reuse, which is different from existing approaches. Specifically, compared with SimFix and ssFix, our approach leverages finer-grained token-level code patterns while SimFix and ssFix cannot as they depend on the AST difference between buggy and referenced code. Similarly, our approach is also different from CapGen as Repatt generates patches by mining code patterns while CapGen only considers the frequency of tokens without considering their combinations. In addition, compared with SearchRepair, our approach does not rely on the heavy semantic code search process. In particular, our approach can leverage nonconsecutive token usage patterns for patch generation, which has been proved to be effective but is not supported by these existing approaches. Additionally, PAR (Kim et al., 2013) and TBar (Liu et al., 2019) aim to improve the patch quality via pre-defining a set of repair templates. Apart from those, a set of APR techniques leverage the advance of mature constraint-solving techniques for patch generation (Nguyen et al., 2013; Mechtaev et al., 2015, 2016, 2018; Gao et al., 2021; Xuan et al., 2017; Chen et al., 2017; Hua et al., 2018). However, since these approaches depend on symbolic execution and constraint solving, they usually face the scalability issue for practical use. Different from these approaches, this work aims to improve the performance of redundancy-based APR techniques and does not face the scalability issue.
With the rapid development of deep learning techniques, the latest APR approaches also leverage such techniques for patch generation. Early studies use statistical machine learning algorithms to sort or pick patch ingredients. For example, Prophet (Long and Rinard, 2016), ACS (Xiong et al., 2017), Elixir (Saha et al., 2017), Hanabi (Xiong and Wang, 2022), and LIANA (Chen et al., 2022) all use different learning models for patch ingredient selection. More recently, state-of-the-art deep learning techniques have been employed in APR techniques, such as SEQUENCER (Chen et al., 2019), CoCoNut (Lutellier et al., 2020), CURE (Jiang et al., 2021), DLFix (Li et al., 2020), Recoder (Zhu et al., 2021), RewardRepair (Ye et al., 2022a), SelfAPR (Ye et al., 2022b), and many others (Fan et al., 2023; Xia et al., 2023). These approaches suffer from interpretability issues and mostly can repair simple bugs (e.g., single-line bugs). In contrast, our approach generates patches by reusing existing code, which complements them based on our experimental results.
8. Conclusion
Redundancy-based APR techniques have been widely studied and many approaches have been proposed. However, they often generate a large number of incorrect patches while only effectively repairing a small number of the bugs. To enhance the performance of redundancy-based APRs, this paper introduces a novel repair technique called Repatt, which employs a two-level pattern mining process to facilitate more precise patch generation through fine-grained code reuse. Evaluations conducted on Defects4J v1.2 and v2.0 benchmarks demonstrate that Repatt complements existing approaches by effectively repairing many unique bugs and achieving higher patch precision in real-world repair scenarios. Furthermore, we present the first attempt at combining different approaches to improve the SoTA APRs. The promising results obtained from this study encourage further research in this direction.